OpenStudy (eric_d):

Factorise a) 10 factorial + 7 factorial

OpenStudy (eric_d):

@ganeshie8

ganeshie8 (ganeshie8):

where are you stuck ?

OpenStudy (eric_d):

frm the beginning hw shld i start

ganeshie8 (ganeshie8):

whats the definition of `factorial` of a number ?

OpenStudy (eric_d):

Product..

ganeshie8 (ganeshie8):

in some sense, `factorial` is just a lazy way of writing product of consecutive integers

ganeshie8 (ganeshie8):

\[\large 5!\] is same as : \[\large 5\times 4\times 3\times 2\times 1\]

OpenStudy (eric_d):

ok

ganeshie8 (ganeshie8):

\[\large 3!\] is same as : \[\large 3\times 2\times 1\]

ganeshie8 (ganeshie8):

can you expand \(10!\) based on above patterns ?

OpenStudy (ikram002p):

mmm why we need to factorize something already simple O.O ?!

OpenStudy (eric_d):

10x9x8x7x6x5x4x3x2x1

ganeshie8 (ganeshie8):

correct! expand \(\large 7!\) also

OpenStudy (eric_d):

7x6x5x4x3x2x1

ganeshie8 (ganeshie8):

guess the teacher wants to give him some practice on factorials @ikram002p :)

OpenStudy (eric_d):

I'm just doing revision..

OpenStudy (anonymous):

[(10)(9)(8)+1](7!)

ganeshie8 (ganeshie8):

@Eyth has it ! 10! = 10x9x8x7x6x5x4x3x2x1 7! = 7x6x5x4x3x2x1 see any common stuff between 10! and 7! ?

ganeshie8 (ganeshie8):

10! = 10x9x8x `7x6x5x4x3x2x1` 7! = ` 7x6x5x4x3x2x1`

OpenStudy (eric_d):

ok..

OpenStudy (anonymous):

Thanks hahaha :) Youre great!

ganeshie8 (ganeshie8):

common stuff is just \(7!\) right ?

OpenStudy (eric_d):

yes

ganeshie8 (ganeshie8):

10! = 10x9x8x `7x6x5x4x3x2x1` 7! = `7x6x5x4x3x2x1` can be writen as : 10! = 10x9x8x `7!` 7! = `7!`

ganeshie8 (ganeshie8):

add them both and factor that common stuff : 10! = 10x9x8x `7!` 7! = `7!` -------------------------- 10! + 7! = ?

OpenStudy (eric_d):

the given answer is 11(9factorial)

ganeshie8 (ganeshie8):

ohk... then i guess we're only half done : 10! = 10x9x8x `7!` 7! = `7!` -------------------------- 10! + 7! = 10x9x8x `7!` + `7!` = `7!` (10x9x8 + 1)

ganeshie8 (ganeshie8):

does that look okay, so far ?

OpenStudy (eric_d):

ys

OpenStudy (eric_d):

what's nxt

ganeshie8 (ganeshie8):

thinking...

ganeshie8 (ganeshie8):

Alright, there seems to be a typo in the given expression, it should be : a) 10 factorial + \(\color{red}{9}\) factorial can you check the question again ?

OpenStudy (ikram002p):

yeah it cant be 11( 9! ) mm

OpenStudy (eric_d):

I had re-check.. There is'nt any typo

OpenStudy (eric_d):

Try this @xapproachesinfinity

OpenStudy (xapproachesinfinity):

So far i found only what ganesh and eyth found! if there no typo that means we need to rewrite this 7!(721)

OpenStudy (eric_d):

Okay @xapproachesinfinity

OpenStudy (xapproachesinfinity):

let me try and see

OpenStudy (xapproachesinfinity):

7!(6!+1) nothing you can achieve here. there must be a typo!

OpenStudy (anonymous):

ganeshie8 is right there must be a typo

OpenStudy (xapproachesinfinity):

Yeah no doubt!